OpenStudy (anonymous):

For a 70-lb force, and a 60-lb force, The resultant force is 100lb. Are the forces pulling at right angles to each other? Show work please

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

this is not at right angle, if they are its supposed to be perfect square Fr=sqrt(70^2+60^2)=100 but they are not

OpenStudy (anonymous):

i can solve the real answer to this prob but the question asked doesnot required it

OpenStudy (anonymous):

 \[F_1+F_2 = 100 \] F1 and F2 are vectors. \[F_1=70i\] \[F_2 = (60cos\theta)i + (60sin\theta)j\] \[70i + (60cos\theta)i+(60sin\theta)j = R\] \[(70 + 60cos\theta)i+(60sin\theta)j = R\] \[\left| R \right| = \sqrt{(70 + 60cos\theta)^2+(60sin\theta)^2}\] \[100 = \sqrt{(70 + 60cos\theta)^2+(60sin\theta)^2}\] \[100^2 = (70 + 60cos\theta)^2+(60sin\theta)^2\] \[100^2-70^2 = 8400cos\theta + 3600cos^2\theta+3600sin^2\theta\] \[17/12 = 7/3cos\theta + cos^2\theta+sin^2\theta\] \[cos^2\theta+sin^2\theta = 1\] \[17/12 = 7/3cos\theta + 1\] \[3(17/12 - 1) /7= cos\theta\] \[\theta = cos^{-1}\left ( 3(17\div12)-1)/7\right )\] \[\theta=79.71^o\] there could be some other easier way to solve, or maybe i am totally wrong :P check answer